OpenStudy (anonymous):

how do you write the equation in slope-intercept form of the line that has a slope of 3 and contains the point (4, 9). HELP ME PLEASE!!!!!!!

OpenStudy (anonymous):

\[y-9=3(x-4)\] is the first step

OpenStudy (anonymous):

i dont understand how you got it...with the points (4,9)

OpenStudy (anonymous):

"point slope" formula \[y-y_1=m(x-x_1)\] a direct substitution with \[x_1=4,y_1=9,m=3\]

OpenStudy (anonymous):

still dont understand

OpenStudy (anonymous):

slope-intercept form is y=mx+b where m is the slope and b is the y-intercept